OpenStudy (studyingallthetime):

write an explicit formula for each recursive formula A(n) = A(n-1)+3.4; A(1) = 7.3 is the explicit formula written A=(7.3-1)+3.4?

OpenStudy (anonymous):

d=A(n)-A(n-1)=3.4 A(1)=7.3 an=a1+(n-1)d A(n)=7.3+(n-1)*3.4

OpenStudy (studyingallthetime):

Thank you very much surjithayer, you help is greatly appreciated

OpenStudy (anonymous):

yw